# Conditional gene genealogies given the population pedigree for a diploid Moran model with selfing

### v1

@@ Abstract
We introduce a stochastic model of a population with overlapping generations and arbitrary levels of self-fertilization versus outcrossing. We study how the global graph of reproductive relationships, or pedigree, influences the genealogical relationships of a sample of two gene copies at a genetic locus. Specifically, we consider a diploid Moran model with constant population size $N$ over time, in which a proportion of offspring are produced by selfing. We show that the conditional distribution of the pairwise coalescence time at a single locus given the random pedigree converges to a limit law as $N$ tends to infinity. This limit law generally differs from the corresponding traditional result obtained by averaging over all possible pedigrees. We describe three different behaviors in the limit depending on the relative strengths, from large to small, of selfing versus outcrossing: partial selfing, limited outcrossing, and negligible outcrossing. In the case of partial selfing, coalescence times are related to the Kingman coalescent, similar to what is found in traditional analyses. In the the case of limited outcrossing, the retained pedigree information forms a random graph coming from a fragmentation-coagulation process, with the limiting coalescence times given by the meeting times of random walks on this graph. In the case of negligible outcrossing, which represents complete or nearly complete selfing, coalescence times are determined entirely by the fixed times to common ancestry of diploid individuals in the pedigree.
@@ 1 Introduction
Modes of reproduction and genetic transmission are fundamental aspects of evolution (Charlesworth and Wright, 2001; Glémin et al., 2019). In this work, we consider autogamy or self-fertilization—hereafter just ‘selfing’—in which meiosis occurs then a zygote is formed by the union of two gametes from the same individual. This form of inbreeding has been especially important in the evolution of plants (Jain, 1976; Uyenoyama et al., 1993; Charlesworth, 2006; Olsen et al., 2021). The common alternative to selfing in plants and animals is outcrossing, in which zygotes are formed by the union of two gametes from different individuals. The selfing rate, defined here as the probability that an individual is produced by selfing, ranges from 00 to $1$. We describe three distinct kinds of coalescent models which hold for different magnitudes of the selfing rate when the process of coalescence is conditioned on the population pedigree, or the organismal genealogy of the population (Ball et al., 1990). This important latent variable has been mostly ignored in theoretical population genetics; see Section 1.3 of this paper and Diamantidis et al. (2024).
@@
The conditional coalescent models we describe exist in the limit as the population size $N$ tends to infinity in the same way that the Wright-Fisher diffusion and the corresponding standard neutral coalescent model exist in this limit (Kingman, 1982; Möhle, 1999; Ewens, 2004). We use $\alpha_{N}$ to denote the selfing rate, with explicit dependence on $N$ so we can control the relative importance of selfing in the model. Our results for coalescence in large populations conditional on the pedigree depend strongly on how $\alpha_{N}$ changes as $N\to\infty$. Most previous analyses, in addition to not conditioning on the pedigree, have assumed a fixed selfing rate. This rate is often denoted $s$, though various notations have been adopted; e.g. $l$ in Haldane (1924), $h$ in Wright (1951), $\beta$ in Pollak (1987) and $\alpha$ in Karlin (1968a, b). As Uyenoyama (2024) recently emphasized, some authors use $s$ for the entire selfing rate while others use it to measure selfing in excess of a background rate $1/N$ due to random mating. Our model includes these as special cases, namely if we assume $\alpha_{N}=s$ or $\alpha_{N}=s+(1-s)/N$, respectively. These two models with fixed $s$ make the same predictions in the limit and both have been called ‘partial self-fertilization’ or ‘partial selfing’.
@@
We use $\alpha\in[0,1]$ to denote the limit of $\alpha_{N}$ as $N\to\infty$. Conditioning coalescence on the pedigree reveals three markedly different behaviors near the boundary $\alpha=1$, depending on how quickly $\alpha_{N}\to 1$. The critical case, which we call ‘limited outcrossing’, is when $1-\alpha_{N}$ is of order $1/N$. Then the conditional coalescent process is characterized by a system of coalescing random walks on a random ancestral graph like the ones previously described for recombination (Griffiths, 1991; Griffiths and Marjoram, 1997) and selection (Krone and Neuhauser, 1997; Neuhauser and Krone, 1997). If outcrossing is even less frequent, for example $1-\alpha_{N}$ of order $1/N^{2}$, the limiting conditional coalescent process is fully determined by the random pedigree of the population, which becomes just a Kingman coalescent tree connecting diploid individuals. We call this ‘negligible outcrossing’. If outcrossing is more frequent than in the critical case, for example $1-\alpha_{N}$ of order $1/\sqrt{N}$, the limiting conditional coalescent model resembles the previously described (unconditional) coalescent model of Nordborg and Donnelly (1997) and Möhle (1998). Here we adopt the previous term ‘partial selfing’. The latter model holds for all $\alpha\in[0,1]$ as long as the rate of approach to $\alpha=1$ is not too fast, whereas limited outcrossing and negligible outcrossing hold just for $\alpha=1$.
@@
In the next two sections, we review previous works which generally use $s$ for the fixed selfing rate, take $s\in[0,1]$ and do not condition on the population pedigree. We will reserve our notation $\alpha_{N}$ for Section 1.3 and later. In Section 1.3, we explain how previous population-genetic models in fact average over pedigrees. In Section 1.1 and Section 1.2 we use $s$ to maintain continuity with what one finds in much of the evolutionary and population-genetic literature.
1.1 Prevalence, causes and consequences of selfing in evolution
@@
Although the ability to self-fertilize is rare among animals (Jarne and Auld, 2006; Avise and Mank, 2009; Escobar et al., 2011) it is found in other taxa including eukaryotic microbes and marine invertebrates (Sasson and Ryan, 2017; Yadav et al., 2023) and is common in plants (Abbott and Gomes, 1989; Hartfield et al., 2017; Teterina et al., 2023). About $10$-$15$% of plant species are predominantly selfing (Wright et al., 2013) as are two of the best-studied genetic model species, Arabidopsis thaliana and Caenorhabditis elegans (Abbott and Gomes, 1989; Sellinger et al., 2020; Barrière and Félix, 2005). The empirical distribution of selfing rates among plant species is markedly bimodal, with relatively fewer species having $s\in(0.2,0.8)$ (Schemske and Lande, 1985; Vogler and Kalisz, 2001; Goodwillie et al., 2005). Genetical models of inbreeding depression can explain this bimodality (Lande and Schemske, 1985; Charlesworth and Willis, 2009). But, whether selfing is favored over outcrossing or vice versa in a given species depends on a number of ecological and evolutionary factors (Charlesworth, 2006; Wright et al., 2013).
@@
Selfing is a form of inbreeding which has three important consequences for the distribution of genetic variation. First, selfing increases single-locus homozygosity. Haldane (1924) described this and the associated decrease in heterozygosity compared to Hardy-Weinberg expectations in an infinite population with selfing rate $s\in[0,1]$. His result may best be described using single-locus probabilities of identity by descent. See Rousset (2002) for a synthetic treatment of identity by descent and other measures of inbreeding and relatedness. In terms of Wright’s individual inbreeding coefficient $F$ (Wright, 1931, 1951), what Haldane (1924) showed was that the reduction in the proportion of heterozygous individuals in the population at equilibrium is given by $F=s/(2-s)$. A simple coalescent interpretation is that $F$ is the probability the two alleles or gene copies at a locus in an individual coalesce in the recent ancestry of the individual, rather than being separated into different individuals by an outcrossing event (Nordborg and Donnelly, 1997).
@@
Second, selfing reduces the effective rate of recombination. Bennett and Binet (1956) discovered this effect by modeling genotype frequencies at two linked loci, and related it to the decrease in single-locus heterozygosity due to selfing. Weir and Cockerham (1973) confirmed this effect in their extensive study of two-locus identity coefficients. Recombination is similarly reduced in finite populations (Golding and Strobeck, 1980; Vitalis and Couvet, 2001). Nordborg (2000) provided an interpretation in terms of coalescence. Backward in time, a recombination event places ancestral genetic material onto two chromosomes. These are necessarily in the same parental individual which might itself have been produced by selfing. Then the same coefficient $F$ is the probability the recombination event gets healed or repaired by a coalescent event within the individual’s recent ancestry. As a result, only a fraction $1-F=2(1-s)/(2-s)$ of recombination events are observable.
@@
Third, selfing generates non-random associations of diploid genotypes at different loci, even if the loci are unlinked. As Smith (2023) recently emphasized, both multi-locus homozygotes and multi-locus heterozygotes are overrepresented. Cockerham and Weir (1968) and Weir and Cockerham (1969, 1973) introduced the term ‘identity disequilibrium’ for this effect, treating it in further detail and with helpful overviews in Cockerham and Weir (1977, 1983, 1984). Haldane (1949) was the first to describe the effect, for inbreeding in general not just selfing, specifically as a consequence of variation in the level of inbreeding among individuals in the population. Bennett and Binet (1956) and Kimura (1958) found and quantified identity disequilibrium for a pair of loci subject to recombination and partial selfing. Weir and Cockerham (1973) partitioned their result into a sum of two terms, one due to linkage, which is equal to zero for unlinked loci, and one due to selfing, which is equal to the theoretical variance of $F$ among individuals in the population:
@@
Weir and Cockerham (1973) considered a number of different coefficients of relatedness, and use $\sigma^{2}_{F_{1}}$ in their version of (1). Narain (1966) described many of the same phenomena as Weir and Cockerham (1973) but in terms of two-locus probabilities of identity by descent directly, rather than in terms of identity disequilibria which are the corresponding (co)variances.
A coalescent interpretation of the identity disequilibrium for two unlinked loci is as follows. Let $Y_{i}$ be an indicator variable, equal to $1$ if the gene copies at locus $i$ in a particular individual coalesce in its recent ancestry, and equal to 00 otherwise. The recent ancestry of a randomly sampled individual is given by the number of generations of selfing in its ancestral line before the first outcrossing event. As Weir and Cockerham (1973) note, this formulation originates with Cockerham and Rawlings (1967). For an individual with $k$ generations of selfing in its immediate ancestry, the probability the two gene copies at a locus remain distinct (i.e. do not coalesce) in any of these generations is $2^{-k}$. If $U$ denotes the random number of selfing generations for an individual,
@@
$\mathrm{Cov}{\left[Y_{i},Y_{j}\right]}=\sum_{k=0}^{\infty}\left(1-2^{-k}\right)^{2}s^{k}(1-s)-\left(\frac{s}{2-s}\right)^{2}$ (4)
which is identical to (1), since $\sigma^{2}_{F}\coloneq\mathrm{Var}{\left[Y_{i}\right]}$ is also given by (4). Interpreting identity disequilibria as covariances is sensible because they occur between loci. In anticipation of later results conditional on the population pedigree, it is important that (2) here is intended as the theoretical distribution for a single individual sampled at random from a given population. In keeping with Haldane (1949), (4) is an average which quantifies variation among individuals.
@@
Owing to these three effects—increased homozygosity, reduced recombination, and identity disequilibrium—partially selfing populations follow different evolutionary paths than randomly mating populations. Selfing alters the evolution of the recombination rate (Charlesworth et al., 1977; Roze and Lenormand, 2005; Stetsenko and Roze, 2022), leads to apparent or spurious multi-locus signals of selection (Cockerham and Rawlings, 1967; Ohta and Cockerham, 1974; Charlesworth, 1991), determines levels of inbreeding depression (Nei, 1965; Hansche et al., 1966; Campbell, 1986; Charlesworth et al., 1990; Lande et al., 1994; Schultz and Willis, 1995; Abu Awad and Roze, 2018), and in many ways alters the responses of populations to selection (Kondrashov, 1985; Wright and Cockerham, 1985; Charlesworth et al., 1993; Nordborg et al., 1996; Willis, 1996; Kelly, 1999a, b, 2007; Lande and Porcher, 2015; Roze, 2016; Sachdeva, 2019). Among these evolutionary studies, Cockerham and Rawlings (1967), Campbell (1986), Kelly (1999a, b, 2007) and Lande and Porcher (2015) explicitly model the variation in the number of selfing generations in the recent ancestry of individuals, given here under neutrality by (2).
The effects of selfing on the distribution of genetic variation are also important determinants of the evolution of the selfing rate (Charlesworth et al., 1990; Uyenoyama and Waller, 1991a, b, c; Abu Awad and Roze, 2020) but in themselves do not explain the prevalence of selfing. In the absence of other evolutionary forces, selfing is strongly favored over outcrossing (Fisher, 1941; Moran, 1962; Karlin and McGregor, 1974; Nagylaki, 1976; Lloyd, 1979; Wells, 1979). As Maynard Smith (1971) observed in discussing the evolution of sexual reproduction or meiosis, outcrossing individuals are at a disadvantage because they transmit only half of their genome to each offspring. The comparison in Maynard Smith (1971) was to parthenogenesis or asexual reproduction, but selfing is a way to alleviate this “cost of sex” in organisms which necessarily undergo meiosis. Thus, while Williams and Mitton (1973) attributed this cost to meiosis itself, Uyenoyama (1986) more precisely reckoned it a cost of outcrossing. The fact that outcrossing is so common can be explained in genetical terms by the avoidance of inbreeding depression (Charlesworth and Willis, 2009; Brown and Kelly, 2020) or the purging or deleterious mutations in the presence of recombination (Kondrashov, 1985; Keightley and Otto, 2006; Kamran-Disfani and Agrawal, 2014).
In addition, many ecological factors affect the evolutionary dynamics of selfing species (Jain, 1976; Solbrig, 1976). Pioneering work by Baker (1955) and Stebbins (1957) showed that selfing may be favored when the usual modes of cross-fertilization are blocked, in plants for example by excessive rainfall or lack of pollinators, or when the environment presents opportunities for colonization starting from one or a few individuals, for example if rare long-distance dispersal events can access open habitat or if there is rapid turnover of suitable habitat patches. Ingvarsson (2002) showed that the last of these, specifically extinction-recolonization dynamics in a metapopulation, could explain the common observation that levels of genetic variation in selfing species are often far below what is predicted by simple neutral population-genetic models (see Section 1.2). Selfing species with their characteristic suite of phenotypes have evolved many times from outcrossing ancestors, prompting questions about whether it might be an evolutionary “dead end” (Stebbins, 1957; Ornduff, 1969; Sicard and Lenhard, 2011; Wright et al., 2013; Cutter, 2019). Within a single species, subpopulations of selfers are commonly restricted to patches at the margins of the habitat (Baker, 1955; Stebbins, 1957), as for example in the case of Leavenworthia alabamica detailed in Busch (2005).
1.2 Population genetics of selfing under neutrality

### v2

@@ Abstract
We introduce a stochastic model of a population with overlapping generations and arbitrary levels of self-fertilization versus outcrossing. We study how the global graph of reproductive relationships, or population pedigree, influences the genealogical relationships of a sample of two gene copies at a genetic locus. Specifically, we consider a diploid Moran model with constant population size $N$ over time, in which a proportion of offspring are produced by selfing. We show that the conditional distribution of the pairwise coalescence time at a single locus given the random pedigree converges to a limit law as $N$ tends to infinity. The distribution of coalescence times obtained in this way predicts variation among unlinked loci in a sample of individuals. Traditional coalescent analyses implicitly average over pedigrees and generally make different predictions. We describe three different behaviors in the limit depending on the relative strengths, from large to small, of selfing versus outcrossing: partial selfing, limited outcrossing, and negligible outcrossing. In the case of partial selfing, coalescence times are related to the Kingman coalescent, similar to what is found in traditional analyses. In the case of limited outcrossing, the retained pedigree information forms a random graph, with coalescence times given by the meeting times of random walks on this graph. In the case of negligible outcrossing, which represents complete or nearly complete selfing, coalescence times are determined entirely by the fixed times to common ancestry of diploid individuals in the pedigree.
@@ 1 Introduction
The organismal genealogy of all individuals in a population for all time, also known as the population pedigree, has not traditionally been of concern in population genetics. But simulation studies, beginning with Ball et al. (1990) and including Kuo and Avise (2008), Wakeley et al. (2016) and Wilton et al. (2017), have suggested that the population pedigree can be important in structuring genetic variation. Of course, some of the earliest work in population genetics concerned how pedigrees affect patterns of variation, specifically identity by descent (Wright, 1921a, b, c, d, e, 1922). But the pedigrees in these early works were either ones laid down by regular systems of mating—see Lachance (2009) for a recent treatment—or pedigrees of arbitrary structure extending back a limited number of generations. These latter, relatively small pedigrees were assumed to be embedded in an infinite population with global properties, such as Hardy-Weinberg equilibrium and no identity by descent, and within which there was no explicit accounting of individuals or their relationships.
Finite population models including coalescent models are models of identity by descent. But population pedigrees did not figure in their development. The approach instead was to average over pedigrees by averaging over the outcomes of reproduction in each generation. See Diamantidis et al. (2024) for more about this history and a discussion of the roles of pedigrees in forward-time versus backward-time models. Averaging over pedigrees is usually done implicitly, as in the statement that two ancestral lines coalesce with probability $1/(2N)$ each generation under the diploid monoecious Wright-Fisher model. It is done explicitly, for example, in equation (4.1) in Kingman (1982) for the coalescence probability under Cannings’ haploid exchangeable model (Cannings, 1974). Predictions for diploid organisms based on this traditional approach are marginal to the population pedigree.
Coalescent theory describes the ancestry of gene copies or alleles in a sample of individuals from a population. Coalescent models are often applied by comparing predicted distributions of genetic variation at a single locus to data from multiple unlinked loci in a sample of individuals. This is done for example in analyses of the site-frequency spectrum (Tajima, 1989; Braverman et al., 1995; Fu, 1995; Achaz, 2009; Gutenkunst et al., 2009; Gao and Keinan, 2016) or when likelihoods are multiplied across unlinked loci (Watterson, 1985; Wakeley, 1999; Nielsen, 2000; Adams and Hudson, 2004; Wang and Hey, 2010; Lohse et al., 2011; Costa and Wilkinson-Herbots, 2017, 2021). But a sample of individuals has a single shared pedigree, fixed in time by past outcomes of reproduction. Further, in large populations with substantial outcrossing, the pedigree of any sample of individuals quickly spreads out to encompass the whole population (Chang, 1999; Derrida et al., 1999; Lachance, 2009; Barton and Etheridge, 2011). So it is important to know whether any specific features of the population pedigree might affect our predictions about genetic variation.
Coalescent models conditional on the pedigree give the correct sampling structure for multiple loci in a sample of individuals. In particular, unlinked loci represent conditionally independent realizations of the process of genetic transmission within the same fixed pedigree. Methods such as multiplying likelihoods across unlinked loci are obviously justified when conditioning on the pedigree. Modeling unlinked loci as independent draws from the distributions of traditionally formulated coalescent models implicitly assumes that each locus has its own population pedigree.
So far there have been only two analytical treatments of coalescence conditional on the population pedigree. Tyukin (2015) proved that the standard neutral coalescent model (Kingman, 1982; Hudson, 1983; Tajima, 1983) can still be used when coalescence is conditioned on the pedigree, for populations which would be described by that model under the traditional approach. But it is not true in general that extensions of the standard neutral coalescent model are similarly applicable to coalescent processes conditional on pedigrees generated under the corresponding population models. Diamantidis et al. (2024) proved this for populations in which very large families occur with some frequency, and went on to describe the resulting conditional coalescent models, in which predictions about genetic variation depend on the fixed times and sizes of large families.
@@
In this work, we consider how coalescence conditional on the population pedigree is affected by autogamy or self-fertilization (hereafter just ‘selfing’). Modes of reproduction are of course fundamental aspects of evolution (Charlesworth and Wright, 2001; Glémin et al., 2019). Selfing is when two gametes produced by meiosis by the same individual unite to form a zygote. This type of inbreeding has been especially important in the evolution of plants (Jain, 1976; Uyenoyama et al., 1993; Charlesworth, 2006; Olsen et al., 2021). The common alternative to selfing in both plants and animals is outcrossing, in which zygotes are formed by the union of two gametes from different individuals. We consider just these two possibilities, and characterize populations by the probability individuals are produced by selfing. We describe three distinct kinds of conditional coalescent models which hold for different magnitudes of the selfing probability.
@@
These models exist in the limit as the population size $N$ tends to infinity in the same way that the Wright-Fisher diffusion and the corresponding standard neutral coalescent model exist in this limit (Ewens, 2004). We use $\alpha_{N}$ to denote the selfing probability, with explicit dependence on $N$ so we can control the importance of selfing in the model. Our results for coalescence in large populations conditional on the pedigree depend strongly on the relative magnitude of $\alpha_{N}$, specifically how $\alpha_{N}$ changes as $N\to\infty$. Most previous analyses, in addition to not conditioning on the pedigree, have assumed a fixed selfing probability. This probability is often denoted $s$, though various notations have been adopted; e.g. $l$ in Haldane (1924), $h$ in Wright (1951), $\beta$ in Pollak (1987) and $\alpha$ in Karlin (1968a, b). As Uyenoyama (2024) recently emphasized, some authors use $s$ for the entire selfing probability while others use it to measure selfing in excess of a background probability of $1/N$ due to random mating. Our model includes these as special cases, namely if $\alpha_{N}=s$ or $\alpha_{N}=s+(1-s)/N$, respectively. Both of these fixed-$s$ models make the same predictions in the limit and both have been called ‘partial self-fertilization’ or ‘partial selfing’.
@@
We use $\alpha\in[0,1]$ to denote the limit of $\alpha_{N}$ as $N\to\infty$. Conditioning coalescence on the pedigree reveals three markedly different behaviors near the boundary $\alpha=1$, depending on how quickly $\alpha_{N}\to 1$. The critical case, which we call ‘limited outcrossing’, is when $1-\alpha_{N}$ is of order $1/N$. Then the conditional coalescent process is characterized by a system of coalescing random walks on a random ancestral graph like the ones previously described for recombination (Griffiths, 1991; Griffiths and Marjoram, 1997) and selection (Krone and Neuhauser, 1997; Neuhauser and Krone, 1997). If outcrossing is even less frequent, for example $1-\alpha_{N}$ of order $1/N^{2}$, the limiting conditional coalescent process is fully determined by the random pedigree of the population, which becomes just a Kingman coalescent tree connecting diploid individuals. We call this ‘negligible outcrossing’. If outcrossing is more frequent than in the critical case, for example $1-\alpha_{N}$ of order $1/\sqrt{N}$, the limiting conditional coalescent model resembles the pedigree-averaged or unconditional coalescent model of Nordborg and Donnelly (1997) and Möhle (1998). Here we adopt the previous term ‘partial selfing’. The latter model holds for all $\alpha\in[0,1]$ as long as the rate of approach to $\alpha=1$ is not too fast, whereas limited outcrossing and negligible outcrossing hold just for $\alpha_{N}\to 1$.
@@
In the next two sections, we review previous works which generally use $s$ for the fixed selfing probability, take $s\in[0,1]$, and do not condition on the population pedigree. We will reserve our notation $\alpha_{N}$ for Section 1.3 and later. In Section 1.3, we explain how previous population-genetic models of selfing in fact average over pedigrees. In Sections 1.1 and 1.2 we use $s$ for continuity with what one finds in much of the evolutionary and population-genetic literature.
1.1 Prevalence and evolutionary-genetic consequences of selfing
@@
Although the ability to self-fertilize is rare among animals (Jarne and Auld, 2006; Avise and Mank, 2009; Escobar et al., 2011) it is found in other taxa including eukaryotic microbes and marine invertebrates (Sasson and Ryan, 2017; Yadav et al., 2023) and is common in plants (Abbott and Gomes, 1989; Hartfield et al., 2017; Teterina et al., 2023). About $10$-$15$% of plant species are predominantly selfing (Wright et al., 2013), as are two of the best-studied genetic model species, Arabidopsis thaliana and Caenorhabditis elegans (Abbott and Gomes, 1989; Sellinger et al., 2020; Barrière and Félix, 2005). The empirical distribution of selfing probabilities among plant species is markedly bimodal, with relatively fewer species having $s\in(0.2,0.8)$ (Schemske and Lande, 1985; Vogler and Kalisz, 2001; Goodwillie et al., 2005). Whether selfing is favored over outcrossing or vice versa in a given species depends on a number of ecological and evolutionary factors (Fisher, 1941; Moran, 1962; Lande and Schemske, 1985; Charlesworth, 2006; Kamran-Disfani and Agrawal, 2014).
@@
Selfing is a form of inbreeding which has three important consequences for the distribution of genetic variation. First, selfing increases single-locus homozygosity via increased identity by descent (Haldane, 1924; Rousset, 2002). In terms of Wright’s individual inbreeding coefficient $F$ (Wright, 1931, 1951), what Haldane (1924) showed was that the reduction in the proportion of heterozygous individuals in the population at equilibrium is given by $F=s/(2-s)$. A simple coalescent interpretation is that $F$ is the probability that the two alleles or gene copies at a locus in an individual coalesce in the recent ancestry of that individual, rather than being separated into different individuals by an outcrossing event (Nordborg and Donnelly, 1997).
@@
Second, selfing reduces the effective rate of recombination (Bennett and Binet, 1956; Weir and Cockerham, 1973; Golding and Strobeck, 1980; Vitalis and Couvet, 2001). Nordborg (2000) provided an interpretation in terms of coalescence. Backward in time, a recombination event places ancestral genetic material onto two chromosomes. These are necessarily in the same parental individual which might itself have been produced by selfing. Then the same coefficient $F$ is the probability that the recombination event gets healed or repaired by a coalescent event within that parental individual’s recent ancestry. So only a fraction $1-F=2(1-s)/(2-s)$ of recombination events is observable.
@@
Third, selfing generates non-random associations of diploid genotypes at different loci, even if the loci are unlinked (Haldane, 1949; Bennett and Binet, 1956; Kimura, 1958; Narain, 1966; Cockerham and Weir, 1968, 1977, 1983, 1984; Smith, 2023; Uyenoyama, 2024). Weir and Cockerham (1973) partitioned this effect into a sum of two terms, one due to linkage, which is equal to zero for unlinked loci, and one due to selfing, which is called the ‘identity disequilibrium’ and is equal to the theoretical variance of $F$ among individuals in the population:
@@
For reference, the quantity (1) is denoted $\sigma^{2}_{F_{1}}$ in Weir and Cockerham (1973, p. 260).
A coalescent interpretation of identity disequilibrium is as follows. Let $Y_{i}$ be an indicator variable, equal to $1$ if the two gene copies at locus $i$ in a particular individual coalesce in its recent ancestry, and equal to 00 otherwise. The recent ancestry of a randomly sampled individual is given by the number of generations of selfing in its ancestral line before the first outcrossing event. Weir and Cockerham (1973) note that this idea originates with Cockerham and Rawlings (1967). For an individual with $k$ generations of selfing in its immediate ancestry, the probability the two gene copies at a locus remain distinct (i.e. do not coalesce) in any of these generations is $2^{-k}$. If $U$ denotes the random number of selfing generations for an individual,
@@
$\mathrm{Cov}{\left[Y_{i},Y_{j}\right]}=\sum_{k=0}^{\infty}\left(1-2^{-k}\right)^{2}s^{k}(1-s)-\left(\frac{s}{2-s}\right)^{2},$ (4)
and this is identical to (1) because $\sigma^{2}_{F}\coloneq\mathrm{Var}{\left(Y_{i}\right)}$ is also given by (4). Interpreting identity disequilibria as covariances is sensible because they occur between loci.
In anticipation of later results conditional on the population pedigree, note that (2) can be interpreted in two different ways. A given pedigree has fixed values of $k$ for every individual in the population. Following Haldane (1949), Weir and Cockerham (1973) and others, (1), (3) and (4) can be interpreted as averages over individuals in an effectively infinite population. Alternatively, (2) can be seen as the prior distribution for a single individual in a given population. In this case, (1), (3) and (4) are seen as averages over all possible population pedigrees. One reason this distinction is important is that while (1) or (4) may be greater than zero, the covariance of $Y_{i}$ and $Y_{j}$ for two unlinked loci in a single individual conditional on the population pedigree must be equal to zero.
@@
Owing to these three effects—increased homozygosity, reduced recombination, and identity disequilibrium—partially selfing populations follow different evolutionary paths than randomly mating populations (Charlesworth et al., 1990, 1993; Uyenoyama and Waller, 1991a, b, c; Lande et al., 1994; Nordborg et al., 1996; Roze and Lenormand, 2005; Roze, 2016; Abu Awad and Roze, 2020). The evolution of selfing also involves a variety of ecological phenomena and associated morphologies (Baker, 1955; Stebbins, 1957; Ornduff, 1969; Jain, 1976; Cutter, 2019). We note that some but not all evolutionary studies have explicitly modeled variation in the number of selfing generations in the recent ancestries of individuals (Cockerham and Rawlings, 1967; Campbell, 1986; Kelly, 1999a, b, 2007; Lande and Porcher, 2015), given here under neutrality by (2).
1.2 Standard neutral models with a fixed selfing probability but no pedigree